OpenStudy (anonymous):

Simplify the following expression. (3^4)^6 ÷ 3^24

OpenStudy (anonymous):

(3^4)^6 / 3^24= 3^(4*6)/3^24= 3^24/3^24= 1

OpenStudy (idealist10):

\[(3^{4})^{6}\div3^{24}\]